OpenStudy (anonymous):

Is anyone willing to help me with chemistry calculations based on moles and molar mass?

OpenStudy (australopithecus):

ask your question don't ask to ask

OpenStudy (australopithecus):

and I can help you with such questions

OpenStudy (anonymous):

are u able to see that^

OpenStudy (anonymous):

I need help on the chart

OpenStudy (anonymous):

?

OpenStudy (australopithecus):

sorry got distracted which problem are you having problems with?

OpenStudy (australopithecus):

I dont want to do your entire worksheet so first off you need to use the formula mole = grams/molecular mass Molecular mass can be found on the periodic table it has units grams/mole Secondly, \[1 mol = 6.022*10^{23} atoms\] avogadro's number = 6.022*10^{23}

OpenStudy (australopithecus):

to find the molecular mass of a compound such as NaOH you would add the molecular mass of Na + Oxygen + Hydrogen then sub that into the formula I provided you

OpenStudy (australopithecus):

that formula being 1mol = grams/molecular mass

OpenStudy (australopithecus):

to convert something from moles to atoms multiply by \[ \frac{6.022*10^{23}atoms}{1mol}\] to convert something from atoms to moles multiply by \[ \frac{1mol}{6.022*10^{23}atoms}\]

OpenStudy (australopithecus):

hope that helps

OpenStudy (australopithecus):

\[mole = \frac{grams}{molecular mass}\]

OpenStudy (australopithecus):

if you have any questions feel free to ask

OpenStudy (australopithecus):

but I think I covered everything that is found on that worksheet

OpenStudy (australopithecus):

number of particles = atoms

OpenStudy (australopithecus):

sorry I should correct this to find the molecular mass of a compound such as NaOH you would add the molecular mass of sodium + Oxygen + Hydrogen. so 22.99g/mol + 15.9994g/mol + 1.0079g/mol = molecular mass of NaOH then sub that into the formula I provided you, if you want to find grams or moles.

OpenStudy (australopithecus):

also next time post this problem in the chemistry section :)